OpenStudy (anonymous):

hve to find circumference of circle when the lengthof its diameter is 2.1cm

OpenStudy (anonymous):

help me plz last ? :P

OpenStudy (anonymous):

circumference of circle= 2 *pi *radius find radius=diameter/2

OpenStudy (anonymous):

r = d/2

OpenStudy (lgbasallote):

\[C = \pi d\] simple as that

OpenStudy (lgbasallote):

where C = circumference d = diameter

OpenStudy (lgbasallote):

\(\pi\) = 3.14

OpenStudy (anonymous):

can u guyz help me by drawing

OpenStudy (anonymous):

here d = 2.1 so r = 2.1/2 circumfrence = 2 pie r = 2 * 22/7 * 2.1/2 = 22 * 0.3 = 6.6 cm

OpenStudy (anonymous):

r=d/2 r=2.1/2=1.05 radius=1.05

OpenStudy (anonymous):

thnx yahoo

OpenStudy (lgbasallote):